The circle C has centre (-1; 5) and radius 5.

i) Show that the equation of C can be written as x2 + y2 + 2x - 10y + 1 = 0.

ii) Find the equation of the tangent to C at the point A ≡ (-4; 9).

The point B lies on the circle C such that AB is a diameter.

(iii) Find the equation of the circle through the points A; B and O, where O is the origin.


1
Expert's answer
2021-09-27T07:35:56-0400

i)


"(x-(-1))^2+(y-5)^2=(5)^2"

"x^2+2x+1+y^2-10y+25=25"

"x^2+y^2+2x-10y+1=0"

ii)


"x^2+y^2+2x-10y+1=0"

Differentiate both sides with respect to "x" and use the Chain Rule


"2x+2yy'+2-10y'=0"

"y'=\dfrac{x+1}{5-y}"

Point "A = (-4, 9)."


"slope=m=y'(-4)=\dfrac{-4+1}{5-9}=\dfrac{3}{4}"

The equation of the tangent to C at the point "A = (-4, 9)" in point-slope form


"y-9=\dfrac{3}{4}(x-(-4))"

The equation of the tangent to C at the point "A = (-4, 9)" in slope-intercept form


"y=\dfrac{3}{4}x+12"

iii) Point "C(-1,5)" is the middle of "AB"


"-1=\dfrac{-4+x_B}{2}=>x_B=2"

"5=\dfrac{9+y_B}{2}=>y_B=1"

Point "B(2,1)."

"(0-m)^2+(0-n)^2=r^2""(-4-m)^2+(9-n)^2=r^2""(2-m)^2+(1-n)^2=r^2"


"m^2+n^2=r^2""16+8m+81-18n=0""4-4m+1-2n=0"


"m^2+n^2=r^2""107-22n=0""m=-\dfrac{1}{2}n+\dfrac{5}{4}"

"m=-\dfrac{13}{11}, n=\dfrac{107}{22}"

"r^2=(-\dfrac{13}{11})^2+(\dfrac{107}{22})^2=\dfrac{12125}{484}"

The equation of the circle


"(x+\dfrac{13}{11})^2+(y-\dfrac{107}{22})^2=\dfrac{12125}{484}"